rawr101:

25P3 A. 12,755 B. 14,290 C. 13,800 D. 18.250

rawr101:

i will fan who can answer

Vocaloid:

permutation formula \[nPr = \frac{ n! }{ (n-r)! }\] for 25P3, n = 25 and r = 3, so our formula becomes \[nPr = \frac{ 25! }{ (25-3)! }=~?\]